OpenStudy (anonymous):

21. A baseball is hit with a bat such that it leaves the bat at a velocity of 30 m/s. it is caught by the center fielder 80 m away from the batter at the same level as it was hit. Find the angle at which the ball left the bat. There seems to be missing knowns to solve this.

OpenStudy (anonymous):

use formula R=(u^2 sin2x)/g x=31.6 degrees

OpenStudy (anonymous):

The distance between batter and fielder is range of the projectile motion of the ball,hence D=range 80 =(u^sin2\[\theta\])/g =>80=(30*30*sin2\[\theta\])/10 i.e sin2\[\theta\]=8/9 hence find theta....

OpenStudy (anonymous):

yup @Taufique but is there any other method without using formula as memorizing dozens of formulas is strenous to the mind??

OpenStudy (anonymous):

how'd you derive this?

OpenStudy (anonymous):

@sarkar there is basic concept here that first of all you calculate time of flight of ball and multiply it with horizontal velocity because of this velocity is constant in overall motion and simply calculate range by apply this basic formula range=horizontal velocity*time of flight...if any query, say me.

OpenStudy (anonymous):

yup i do understand the derivation of that quite clearly...